2 players has got a total of 286 hits one has 14 more hites than one

Answer by checkley79(3050)

X+(X+14)=286
2X=286-14
2X=272
X-272/2
X=136 ANS. FOR THE LOWEST HITTER.
136+14=150 ANS. FOR THE BETTER HITTER.
PROOF:
136+150=286
286=286